# A Uniform Rewriting Algorithm for Twisted Tensor Representations of Finite General Linear Groups

### v1

@@ 7. Computational experiments
In this section, we report on some small computational experiments which serve as a sanity check for the number-theoretic Lemma 3.1 and for the spectral behaviour of Singer cycles on symmetric powers of the natural module, as well as a toy model for the reconstruction step in the rewriting algorithm. All computations were performed in SageMath; the corresponding script is available at https://github.com/phucdv2018/singer_sym_check.sage/tree/v1.0.1 (archived at DOI: https://doi.org/10.5281/zenodo.17744874) and included as a supplementary file singer_sym_check.sage. The code is intended as a collection of sanity checks for the key mechanisms in the paper and is not optimised for large parameters.
@@ 7.2. Model eigenvalues on tensor powers
$\texttt{Example weight pattern c = (0, 0, 3)},$

### v2

@@ 20 E = sum(b[i] * (q**i) for i in range(d)) % modulus
Note. The implementation above uses itertools.product which is suitable only for small parameters. For the large-scale experiments in Section 7.2 below (e.g., $q=2^{16},d=10$), our supplementary script employs an optimized recursive generator weight_patterns_sumK and performs arithmetic purely on integer exponents modulo $q^{d}-1$, avoiding the expensive construction of the extension field.
@@ 7. Computational experiments
In this section, we report on some small computational experiments which serve as a sanity check for the number-theoretic Lemma 3.1 and for the spectral behaviour of Singer cycles on symmetric powers of the natural module, as well as a toy model for the reconstruction step in the rewriting algorithm. All computations were performed in SageMath; the corresponding script is available at https://github.com/phucdv2018/singer_sym_check.sage (archived at DOI: https://doi.org/10.5281/zenodo.17792033) and included as a supplementary file singer_sym_check.sage. The code is primarily intended as a collection of sanity checks for the key mechanisms in the paper. While the matrix reconstruction routines are designed for small examples to illustrate correctness, the eigenvalue labeling and injectivity checks are implemented efficiently and scale to large parameters.
@@ 7.2. Model eigenvalues on tensor powers
Example weight pattern c = (0, 0, 3),
@@
As a more demanding sanity check closer to the intended applications, we also ran a purely “exponent–model” variant of this experiment for larger parameters. Specifically, we considered
$(q,d,K)\;=\;(2^{16},\,10,\,K)\qquad\text{with}\qquad K\in\{2,3,4\}.$
In this setting we do not construct $\mathbb{F}_{q^{d}}$ or a Singer matrix explicitly, but work only with the exponents $E(\mathbf{c})$ in $\mathbb{Z}/(q^{d}-1)\mathbb{Z}$. For each $K$ we enumerate all digit vectors $\mathbf{c}\in\mathcal{B}_{K}$ of length $d$ with $\sum_{i}c_{i}=K$ and verify that the map
$\mathbf{c}\;\longmapsto\;\sum_{i=1}^{d}c_{i}q^{i-1}\bmod(q^{d}-1)$
is injective. For $K=2,3,4$ we have
$|\mathcal{B}_{K}^{\prime}|\;=\;\binom{d+K-1}{K}\;=\;55,\ 220,\ 715,\ \ \mbox{respectively},$
and in each case the program reports that all $|\mathcal{B}_{K}^{\prime}|$ exponents are distinct and that no collisions occur. This confirms the base-$q$ injectivity lemma in a regime where the module dimension reaches
$\dim\mathrm{Sym}^{4}(V)\;=\;\binom{10+4-1}{4}\;=\;715,$
while avoiding any explicit matrix computations over the large field $\mathbb{F}_{2^{160}}$.
Finally, we demonstrate the feasibility of our labeling strategy for these parameters. Our implementation generates the lookup table for $K=4$ and verifies injectivity modulo $2^{160}-1$ in approximately $30$ milliseconds on a standard workstation. This confirms that, by working directly with exponents and look-up tables rather than discrete logarithms, the eigenvalue identification step has negligible computational cost compared to matrix operations, even over very large fields.
@@ 7.5. Discussion
  • the labeling strategy scales efficiently to large groups such as $SL_{10}(2^{16})$, identifying weights in millisecond time without the need for explicit field arithmetic in $\mathbb{F}_{2^{160}}$, as demonstrated in Section 7.2;